# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v2

@@ 4.2 High-Dimensional Performance Analysis (n=50, k=10)
We now visually examine the performance in the most challenging setting: $n=50$ assets and $k=10$ state variables. This scenario tests scalability where traditional DP methods are intractable. As shown quantitatively in Table 2 and Figure 1c, the 2-PG-DPO method achieves a minimum RMSE ($1.32\times 10^{-2}$ at epoch 3,200) significantly better than the Baseline approach ($1.26\times 10^{-1}$ at epoch 100,000).
The visualizations in Figures 2 and 3 focus on the policy for the first asset ($i=1$) versus time-to-maturity ($T-t$) and the first state variable ($Y_{1}$), holding other state variables constant at their long-term means. The results shown correspond to the state on epoch 3,200, near where the 2-PG-DPO method achieved its minimum observed RMSE.
@@
(b) 2-PG-DPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
(c) Baseline Error
(d) 2-PG-DPO Error
Figure 2: Comparison of total policies and errors for the first asset ($i=1$) in the n=50, k=10 case on epoch 3,200 (plotted against $Y_{1}$). Top row compares policies, bottom row compares errors (using Symlog scale). Note the significantly different color bar ranges are used for each plot to best visualize its content, reflecting the dramatic difference in policy/error magnitudes.
(a) 2-PG-DPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
(b) 2-PG-DPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
@@
Figure 3: Decomposition of the 2-PG-DPO policy for the first asset ($i=1$) in the n=50, k=10 case on epoch 3,200 (plotted against $Y_{1}$). Top row shows myopic and hedging components. Bottom row shows their respective errors (using Symlog scale). Note the use of different color bar ranges, necessary to visualize both the negligible myopic error and the larger hedging error, which dictates the total policy error.
The challenges of high dimensionality become starkly evident for the Baseline PG-DPO method in Figure 2. While the learned policy (2(a)) shows some structure, the corresponding error (2(c)) is extremely large, visually confirming the poor quantitative performance reported in Table 2. This indicates severe inaccuracies in the policy learned by the Baseline method when faced with the combined complexity of 50 assets and 10 state variables.
In dramatic contrast, the 2-PG-DPO method provides a highly accurate solution, as shown by its policy (2(b)) and error (2(d)). Comparing the value ranges indicated by the color bars, the error magnitude for the 2-PG-DPO policy is roughly an order of magnitude smaller than the Baseline error, underscoring the robustness and practical necessity of the 2-PG-DPO approach. Although perfect accuracy might be harder to achieve as dimensionality increases (the minimum RMSE of $1.32\times 10^{-2}$ is higher than in lower-dimensional cases), it remains remarkably effective and vastly superior to the direct policy learning method.
The component decomposition in Figure 3 confirms the patterns observed previously. The myopic component (3(a)) largely dictates the policy shape, and its associated error (3(c)) remains negligible (confirming quantitative findings that this error is often $O(10^{-7})$), requiring a very small scale on its color bar for visualization. The hedging component (3(b)) contributes visibly to the total policy shape. Crucially, the hedging error (3(d)) is visually almost identical in pattern to the total 2-PG-DPO error (2(d)) when comparing their respective plots, confirming that the overall accuracy is dictated by the precision of estimating the hedging demands.
In conclusion, the $n=50,k=10$ results powerfully demonstrate the capability of the 2-PG-DPO framework to “break the dimensional barrier”. While the Baseline PG-DPO struggles significantly, 2-PG-DPO delivers a highly accurate and viable solution by leveraging the structure of Pontryagin’s Maximum Principle. Its performance remains remarkably strong and vastly superior to the baseline, solidifying its potential as a practical tool for large-scale, continuous-time portfolio optimization.
@@ 5 Conclusion
This paper tackled the significant challenge posed by the “curse of dimensionality” in solving large-scale continuous-time portfolio optimization problems. We introduced and developed the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, which uniquely combines the theoretical rigor of Pontryagin’s Maximum Principle (PMP) with the power of deep learning, specifically using backpropagation-through-time (BPTT) applied to policy networks representing investment and consumption strategies.
A central contribution is the development of a highly effective “Two-Stage” PG-DPO variant. This method leverages the insight that BPTT applied to simulation paths can yield stable estimates of the crucial Pontryagin costates (and their derivatives) relatively quickly, often much faster than the full policy network converges. By utilizing these rapidly obtained costate estimates within the analytical first-order optimality conditions derived from PMP, the Two-Stage approach efficiently constructs near-optimal control policies, effectively decoupling the estimation of marginal value sensitivities from the full policy optimization.
Our numerical experiments, conducted on a multi-asset, multi-factor affine model chosen for its tractability allowing rigorous benchmark comparisons, demonstrated the practical effectiveness and scalability of this approach. The results consistently showed that the 2-PG-DPO method achieves substantially higher accuracy and faster convergence compared to the baseline end-to-end policy learning strategy. Notably, the framework successfully scaled to problems involving dimensions (e.g., n=50 assets, k=10 state factors) that are typically considered intractable for traditional numerical HJB/PDE solvers, thereby effectively “breaking the dimensional barrier.” We specifically highlighted the method’s capability to accurately capture both myopic and complex intertemporal hedging demands by explicitly leveraging the structure inherent in the PMP optimality conditions.
While demonstrated on an affine model, the underlying principles relying on PMP and BPTT suggest broader applicability. This research opens several avenues for future investigation. Firstly, further exploration of the theoretical underpinnings is crucial. Our analysis, related to Proposition 3.1, indicates that the rationale behind 2-PG-DPO—leveraging the stability of value function derivatives—is theoretically supported for potentially broad classes of models, possibly including non-affine cases, provided certain regularity conditions related to smoothness and parabolicity are met. A key avenue for future theoretical research is to rigorously characterize the precise range of financial models (both affine and non-affine) that satisfy these sufficient regularity assumptions, thereby ensuring the well-posedness of the PMP framework and the reliability of BPTT-based estimates. Furthermore, investigating the general theoretical convergence properties of the PG-DPO algorithm, particularly the costate estimation process, remains an important open question.
Secondly, extending the framework to incorporate more realistic market features is a natural next step for enhancing practical relevance. This includes adapting PG-DPO to handle portfolio constraints (like short-selling restrictions or leverage limits) and modeling market frictions such as transaction costs, both of which significantly alter the optimal control landscape. Testing the framework’s robustness and performance with different utility functions beyond CRRA (e.g., HARA, prospect theory) is also warranted. Moreover, explicitly including the optimization and evaluation of the intermediate consumption policy, omitted in the current study for benchmarking clarity, would provide a more complete assessment of the full framework’s capabilities.
Thirdly, further empirical validation and algorithmic refinements are desirable. Building on the theoretical potential for broader applicability, applying the PG-DPO framework (especially the superior 2-PG-DPO variant) to well-known non-affine financial models (e.g., stochastic volatility models like Heston) would be valuable, even in the absence of exact benchmarks, to showcase its capabilities more broadly. Algorithmically, exploring alternative neural network architectures, advanced optimization techniques, and variance reduction methods beyond antithetic variates could potentially lead to further performance improvements or faster convergence.
Finally, optimizing the implementation details deserves attention. This includes refining strategies for determining the optimal warm-up duration for the 2-PG-DPO method to best balance costate accuracy against potential overfitting to simulation artifacts inherent in the discrete-time training process. Careful tuning of hyperparameters related to both the network training and the costate estimation phase will be important for practical applications.
@@ Appendix B Analytic Solutions for Affine Models
For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$. Following Kim & Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form:

### v3

@@ 4.2 High-Dimensional Performance Analysis ($n=50,k=10$)
We now visually examine the performance of the three methodologies in the most challenging setting considered: $n=50$ assets and $k=10$ state factors. This high-dimensional scenario, where traditional DP methods are intractable, serves to highlight the distinct capabilities and limitations of each approach. As a reference, Figure 2 displays the benchmark analytical solution for the first asset ($\pi_{1}^{\text{sol}}$), decomposed into its myopic ($\pi_{1,0}^{\text{sol}}$) and intertemporal hedging ($\pi_{1,1}^{\text{sol}}$) components. The hedging component is clearly non-trivial and exhibits complex dependencies on both the state variable $Y_{1}$ and time-to-maturity, underscoring its importance for optimal allocation.
(a) Benchmark Policy ($\pi_{1}^{\text{sol}}$)
(b) Myopic Comp. ($\pi_{1,0}^{\text{sol}}$)
(c) Hedging Comp. ($\pi_{1,1}^{\text{sol}}$)
Figure 2: Benchmark (analytical) solution components for the first asset ($i=1$) in the $n=50,k=10$ case, plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (e.g., $t=(T-T_{0})/2$ if $T_{0}$ is initial time, assuming $T=1.5$ is the horizon from $t=0$; thus, for instance, at $t=0.75$ or time-to-maturity $0.75$). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. From left to right: total optimal policy, its myopic component, and its intertemporal hedging component. These serve as the ground truth.
Figure 3 provides a side-by-side comparison of the policies learned by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and Deep BSDE, alongside their respective errors, for the first asset. These plots correspond to iterations where each method achieved its (near) optimal performance in the $n=50,k=10$ setting, as indicated in Table 2 (e.g., iteration 5400 for P-PGDPO, which is the nearest logged checkpoint to the observed minimum RMSE at iteration 5399, as evaluations were performed every 200 iterations).
@@
(b) Baseline Error
(c) P-PGDPO Policy ($\pi_{1}^{\mathrm{PMP}}$)
(d) P-PGDPO Error
(e) Deep BSDE Policy ($\pi_{1}^{\mathrm{BSDE}}$)
(f) Deep BSDE Error
Figure 3: Comparison of total policies (left column) and policy errors (right column) for the first asset ($i=1$) across Baseline PG-DPO (top row), Projected PG-DPO (P-PGDPO) (middle row), and Deep BSDE (bottom row) in the $n=50,k=10$ case. Results shown at (near) optimal iterations for each method (e.g., iteration 5400 for P-PGDPO), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. Errors use a symmetric logarithmic (symlog) scale. Different color bar ranges reflect varying policy/error magnitudes.
The Baseline PG-DPO method (Figure 3a,b), despite utilizing a standard neural network architecture (detailed in Appendix D.2), struggles significantly. While its learned policy exhibits some non-linear structure, it bears little resemblance to the benchmark, and the corresponding error is extremely large. This performance visually confirms the poor quantitative results in Table 2 and highlights the difficulty of directly approximating the high-dimensional optimal policy function using a neural network without leveraging more structural information from the problem.
(a) P-PGDPO Myopic ($\pi_{1,0}^{\mathrm{PMP}}$)
(b) P-PGDPO Hedging ($\pi_{1,1}^{\mathrm{PMP}}$)
@@
Figure 4: Decomposition of the Projected PG-DPO (P-PGDPO) policy for the first asset ($i=1$) in the $n=50,k=10$ case (results from iteration 5400), plotted against $Y_{1}$ at a representative time-to-maturity $T-t$ (consistent with Figure 2). Other state factors $Y_{j\neq 1}$ are held at their long-term means $\theta_{Y,j}$. Top row: myopic and hedging components. Bottom row: their respective errors (symmetric logarithmic (symlog) scale). Note different color bar ranges.
The Deep BSDE method (Figure 3e,f) shows a marked improvement over the Baseline. Its learned policy qualitatively resembles the myopic component of the benchmark (cf. Figure 2b), primarily capturing dependencies on the state variable $Y_{1}$. However, the policy largely fails to capture the complex intertemporal hedging demands. The error plot for Deep BSDE (Figure 3f) shows significant discrepancies in regions where hedging is important, suggesting that the hedging component is largely absent from the Deep BSDE solution. This difficulty in capturing hedging demands can be attributed to several factors. The Deep BSDE method primarily focuses on matching the terminal condition and ensuring consistency of the value process $Y_{t}$ and the gradient-related process $Z_{t}$ (approximating first-order derivatives of the value function) through its loss function. Intertemporal hedging, however, critically depends on the mixed second-order derivative $V_{x\mathbf{y}}$, which is not explicitly represented or directly targeted by the standard two-network Deep BSDE architecture. Furthermore, the loss function penalizes inaccuracies in $Z_{t}$ primarily through a quadratic martingale term. Consequently, small errors in estimating the gradient components relevant to $V_{x\mathbf{y}}$ might have a limited impact on the overall training objective but can be significantly magnified when constructing the policy, especially after multiplication by the potentially ill-conditioned inverse covariance matrix $\Sigma^{-1}$ in high-dimensional settings. As a result, while the Deep BSDE approach can effectively learn the dominant myopic component of the policy, it may struggle to precisely fit the more nuanced intertemporal hedging patterns.
In stark contrast, the Projected PG-DPO (P-PGDPO) method (Figure 3c,d) delivers a policy that is visually almost identical to the benchmark total policy (Figure 2a). The corresponding error is substantially smaller than that of the other two methods. This superior performance underscores the efficacy of its decoupled learning structure: first estimating costates via BPTT and stabilizing them through Monte Carlo averaging, then analytically constructing the policy using the PMP FOCs.
The remarkable accuracy of P-PGDPO is further elucidated by decomposing its policy into myopic and hedging components, as shown in Figure 4 (results from iteration 5400). The myopic component (Figure 4(a)) is virtually identical to the benchmark myopic component, with its error (Figure 4(c)) being negligible ($O(10^{-7})$). The hedging component derived by P-PGDPO (Figure 4(b)) also successfully captures the complex shape of the benchmark hedging demand (Figure 2c). The remaining error in the hedging component (Figure 4(d)) dictates the overall error of the P-PGDPO policy. This residual hedging error is likely not due to a fundamental flaw in the P-PGDPO algorithm itself but rather a consequence of the numerical challenges inherent in high-dimensional asset allocation. Specifically, the hedging term in the PMP FOC involves $\Sigma^{-1}$:
$\pi^{\text{hedge}}=-\frac{1}{X_{t}V_{xx}}\,\underbrace{\Sigma^{-1}}_{\text{cond num can be large for }n\uparrow}(\bm{\sigma}\,\bm{\rho}\,\bm{\sigma}_{Y})\,V_{x\mathbf{Y}}.$
For $n=50$, the $50\times 50$ covariance matrix $\Sigma$ can easily become ill-conditioned. Even if the BPTT-derived estimate for $V_{x\mathbf{Y}}$ is quite accurate, any small residual noise can be significantly amplified when multiplied by $\Sigma^{-1}$ if its condition number is large. This amplification effect is the primary suspect for the observed hedging error.
In conclusion, the $n=50,k=10$ visualizations powerfully demonstrate that Projected PG-DPO (P-PGDPO) can effectively ”break the dimensional barrier” where Baseline PG-DPO fails and Deep BSDE offers only a partial (myopic-focused) solution. The remaining inaccuracies in the hedging component point towards the need for addressing the ill-conditioning of the covariance matrix in very high asset dimensions.
@@ 5 Conclusion
This paper confronted the enduring “curse of dimensionality” in continuous-time portfolio optimization, a challenge that manifests distinctly with increasing asset numbers ($n$) versus state factors ($k$). Our numerical investigations (Section 4) revealed that while deep learning methods manage larger $k$ relatively well, escalating $n$ poses severe difficulties, particularly due to the ill-conditioning of asset covariance matrices. Furthermore, we identified that prevalent approaches like Deep BSDE, while proficient at capturing myopic policy components, often fail to adequately represent crucial intertemporal hedging demands, a limitation rooted in their network architecture and loss functions which do not explicitly target the necessary second-order mixed derivatives of the value function (e.g., $V_{x\mathbf{y}}$).
To surmount these obstacles, we introduced the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework, with its highly effective Projected PG-DPO (P-PGDPO) variant (Section 3.2). P-PGDPO distinctively leverages Pontryagin’s Maximum Principle (PMP) by first employing backpropagation-through-time (BPTT) to achieve rapidly stabilizing estimates of the Pontryagin costates and their derivatives (e.g., $\lambda_{t},\partial_{x}\lambda_{t},\partial_{\mathbf{Y}}\lambda_{t}$). These estimates are then analytically projected onto the manifold of PMP-optimal controls via the first-order conditions. This decoupling strategy proved remarkably successful, enabling P-PGDPO to tackle problems of previously prohibitive scales. For instance, in a challenging $n=50$ asset, $k=10$ factor environment, P-PGDPO achieved a total policy RMSE of $\mathbf{1.58\times 10^{-2}}$ relative to the analytical benchmark, with the myopic component error being negligible ($O(10^{-7})$) and, critically, accurately reconstructing complex hedging terms (Section 4.2, Figure 4). The theoretical underpinnings of P-PGDPO are strengthened by our Policy Gap Bound (Theorem 1), which guarantees near-optimal policies given accurate costate estimation and a small FOC gap.
Despite its successes, P-PGDPO’s accuracy in representing hedging terms can be affected by the amplification of estimation noise through the inverse covariance matrix $\Sigma^{-1}$ when $n$ is large and $\Sigma$ is ill-conditioned (Section 4.2). Concurrently, our experiments confirmed that Deep BSDE, while adept at matching myopic demands, struggles with learning the nuanced cross-derivative information essential for hedging, largely due to its architectural and objective function design (Figure 3). These observations highlight the distinct strengths and weaknesses of different deep learning paradigms in high-dimensional control.
The practical implications of P-PGDPO are significant. Its demonstrated ability to recover complex, non-linear hedging strategies in high dimensions suggests its potential applicability to large institutional portfolios, such as those involving numerous ETFs or derivative hedging. Moreover, the core principle of P-PGDPO—“estimate costates first, then derive policy”—holds promise for a broader class of problems, where PMP provides the necessary optimality structure (Section 3.3).
Future research will pursue several promising directions:
  • Developing “P-PGDPO 2.0” by integrating $\Sigma$ regularization or shrinkage techniques to mitigate the impact of ill-conditioned covariance matrices on hedging term estimation.
  • Extending the P-PGDPO framework to incorporate portfolio constraints (e.g., no short-selling, leverage limits) and market frictions like transaction costs, by adapting the PMP conditions accordingly.
  • Investigating the performance of P-PGDPO with non-HARA utility functions and behavioral finance objectives, and fully integrating the optimization of intermediate consumption policies ($C_{t}$).
  • Validating the framework through rigorous backtesting on historical and simulated market data for non-affine models.
  • Rigorously characterizing the precise range of financial models that satisfy the sufficient regularity assumptions for the P-PGDPO framework’s effectiveness.
In essence, PG-DPO, particularly its projected variant, demonstrates that the synergy between deep learning and the structural guidance of Pontryagin’s Maximum Principle can redefine the boundaries of solvability for high-dimensional continuous-time financial control problems.
@@ Appendix B Analytic Solutions for Affine Models
For simplicity, we assume no intermediate consumption, i.e., $C_{t}=0$, and consider an investor who derives utility only from terminal wealth $X_{T}$ under a CRRA objective $U(x)=\frac{x^{1-\gamma}}{1-\gamma}$. Following Kim and Omberg, (1996), the associated Hamilton–Jacobi–Bellman (HJB) equation can be solved by positing a value function of the form: